# Generative Adversarial Networks in Ultrasound Imaging: Extending Field of View Beyond Conventional Limits

### v1

@@ Abstract
Transthoracic Echocardiography (TTE) is a fundamental, non-invasive diagnostic tool in cardiovascular medicine, enabling detailed visualization of cardiac structures crucial for diagnosing various heart conditions. Despite its widespread use, TTE ultrasound imaging faces inherent limitations, notably the trade-off between field of view (FoV) and resolution. This paper introduces a novel application of conditional Generative Adversarial Networks (cGANs), specifically designed to extend the FoV in TTE ultrasound imaging while maintaining high resolution. Our proposed cGAN architecture, termed echoGAN, demonstrates the capability to generate realistic anatomical structures through outpainting, effectively broadening the viewable area in medical imaging. This advancement has the potential to enhance both automatic and manual ultrasound navigation, offering a more comprehensive view that could significantly reduce the learning curve associated with ultrasound imaging and aid in more accurate diagnoses. The results confirm that echoGAN reliably reproduce detailed cardiac features, thereby promising a significant step forward in the field of non-invasive cardiac naviagation and diagnostics.
Introduction
Transthoracic Echocardiography (TTE) Ultrasound (USG) is a cornerstone diagnostic tool in cardiovascular medicine, valued for its non-invasive approach and ability to comprehensively assess cardiac anatomy and function. It enables detailed visualization of cardiac structures, such as chambers, valves, and myocardium, and is essential in diagnosing conditions like heart failure, valvular heart disease, and cardiomyopathies. One of TTE USG’s key advantages is its capacity for real-time dynamic imaging, allowing clinicians to observe cardiac motion and provide insights beyond what static imaging can offer.
@@ Introduction
In this context, our research has explored the application of a conditional Generative Adversarial Networks (cGANs) to extend the FoV in TTE USG imaging while maintaining high resolution. The proposed cGAN architecture, denoted as echoGAN, is capable of generating realistic anatomical structures through outpainting, thereby broadening the viewable area in medical imaging. This approach aims to address the limitations in FoV and resolution trade-offs, potentially enhancing both automatic and manual ultrasound navigation. For automatic navigation [1], an expanded FoV can provide a more comprehensive view, potentially improving the efficiency and accuracy of imaging processes, especially in complex diagnostic scenarios.
@@
Furthermore, this extended FoV could be particularly beneficial for less experienced physicians in manual navigation.The study by Chisholm et al. [2] and the work of Vinodth et al. [3] focus on cardiac ultrasound training, specifically evaluating the duration of training necessary to achieve proficiency in TTE. These studies demonstrate that mastering TTE requires a considerable amount of time. By offering a wider yet detailed view, our cGAN-based approach could reduce the learning curve associated with ultrasound imaging and assist in making more accurate diagnoses. This is especially relevant in cases requiring both detailed examination and broader assessments.
Guidance during ultrasound imaging, especially in echocardiography, presents significant challenges. While previous studies, such as the one by Li et al. [4], have explored automated robotic navigation in Transesophageal Echocardiography (TEE) with limited degrees of freedom, similar challenges persist in TTE. The complex nature of cardiac anatomy and the need for precise imaging in both TEE and TTE demand high expertise and skill from echocardiographers [3].
In conventional TEE examinations, the echocardiographer manually guides the probe to acquire standard views of the heart’s anatomy, as outlined in guidelines by Hahn et al. [5]. This process involves intricate navigation, as the echocardiographer must rely on their interpretation of ultrasound images and a mental model of the heart’s spatial layout. Achieving proficiency in this task requires substantial training and experience. Similarly, in TTE, despite the non-invasive nature of the procedure, obtaining optimal imaging planes is a complex skill. The echocardiographer must maneuver the probe on the patient’s chest, often working within limited space and requiring a nuanced understanding of cardiac structures and their orientation. This complexity is heightened by the variability in patients’ anatomy and the need to adjust probe positions to capture clear images of specific cardiac areas.
Both TEE and TTE share the common challenge of probe navigation, heavily reliant on the operator’s skill and understanding of cardiac anatomy. The difficulty lies in aligning the ultrasound beam with the heart’s structures to obtain the necessary views for accurate diagnosis. This process is often not straightforward and demands a high level of dexterity and spatial awareness. Given these challenges, there is a continuous search for technological advancements that can aid in simplifying the navigation process, making echocardiography more accessible and reliable, especially for less experienced practitioners.
@@
In the domain of medical imaging, particularly ultrasound imaging, the task of outpainting introduces a set of unique challenges that are markedly distinct from those encountered in the outpainting of general or natural images [7, 8, 9]. The paramount requirement for anatomical accuracy in medical images necessitates that outpainting algorithms not only generate visually plausible extensions but also ensure these extensions accurately represent the expected anatomy. This is crucial, as any deviation could compromise the clinical utility of the image. Moreover, ultrasound images are characterized by specific textures and patterns that denote various tissues, fluids, and pathological conditions. Ensuring continuity in these textures and patterns during the outpainting process is essential, yet challenging, as the algorithm must replicate medical characteristics that might be absent from the portion of the image initially visible. This is further complicated by the imperative to maintain diagnostic integrity; the outpainted areas must not introduce artifacts or inaccuracies that could lead to misinterpretation or misdiagnosis.
Results
Dataset
The data utilised in this study comprises clinical examinations derived from 500 patients, obtained at the University Hospital of St Etienne in France collected under the CAMUS project [10]. The data acquisitions were aimed to facilitate Left Ventricular Ejection Fraction (LVEF) measurements. For each patient, 2D apical fourchamber and two-chamber view sequences were recorded. The images were extracted from video recordings yielding in average 20 images per patient (min 10, max 42). Alltogether 9964 images constitute the final dataset. From these, 7917 corresponding to 400 patients were used for training of generative adversarial network. To avoid data leakage, it was ensured that data from the same patient are present only in training or only in testing dataset, not both.
Outpainting the Ultrasound FoV
@@ Results
This approach not only promises to refine navigation accuracy but also to facilitate a more comprehensive understanding of the heart’s structure and function. Enhancing the FoV through outpainting addresses a critical gap in cardiac imaging by ensuring that the right heart is not overlooked, thus supporting more informed clinical decision-making. Our findings could mark a significant step forward in non-invasive cardiac diagnostics, offering a novel solution to the longstanding challenge of incomplete cardiac visualization during echocardiography.
Different size of extended view
We evaluated the influence of the training cone size on the quality of outpainted sections in ultrasound images. Figure 3 illustrate two distinct scenarios: the first utilizes a narrower 10-degree cone cut as the inference input, with the echoGAN algorithm tasked with generating outpainted extensions that span 40 degrees on both sides; the second employs a 60-degree cone cut for inference, with outpainting applied to 15-degree segments on each side. Notably, the generated images appear highly realistic, even when echoGAN is interfaced with only a 10-degree cut. The algorithm demonstrates ability to reproduce cardiac muscle structures with high fidelity. This is particularly evident in the first, second, and fourth rows, where echoGAN successfully outpaints a complete right atrium, that is an encouraging result given that the right atrium was not included in the 10-degree cut, suggesting that echoGAN can reconstruct it entirely. In the case of the left ventricle and left atrium, we observe that echoGAN is capable of outpainting the missing boundary parts for the 30-degree cut and can significantly reconstruct substantial portions of these structures for the 10-degree cut. Again, to quantitavely evaluate the quality of outpainted USG images, we measure FID for different sizes of outpainted regions. The results are provided in Tab. 1. The 30 degree cut means, that echoGAN generates 15 degrees on both sides. Similarly in case of 80 degrees, echoGAN geenrates 40 degree on both sides, using only very narrow 10 degree cone as an input.
Cut 30 Cut 46 Cut 60 Cut 80
FID 122.08 143.83 145.74 149.63
Table 1: The average FID for different sizes of outpainted area.
Right ventricle area
Our experiments are specifically aimed at the quantification of the RV area from both original and outpainted echocardiograms. The rationale for focusing on the RV area is influenced by the established practice in cardiology where, for instance, the measurement of left ventricular ejection fraction incorporates the left ventricular area as a crucial parameter in its evaluation [12]. Drawing inspiration from such methodologies, we hypothesize that if the RV areas derived from generated (outpainted) and real echocardiograms are similar, it would substantiate the utility and validity of the generated images for clinical and diagnostic purposes.
We performed the Wilcoxon signed-rank test to compare the RV areas obtained from the original and the outpainted echocardiograms. The results of this test confirmed the null hypothesis, indicating that there is no statistically significant difference between the RV areas measured in the original echocardiographic images and those derived from images extended through echoGAN outpainting. This finding validates the accuracy and effectiveness of the GAN-based outpainting technique in echocardiography. The ability of GANs to generate anatomically coherent extensions of the cardiac structures without compromising the integrity of the images suggests that this approach can reliably extend the field of view in echocardiographic navigation.
Figure 4: Segmentation over Right Ventricle for GT 4CH view with
Discussion
Methods
Image preprocessing
Due to the absence of US data in raw format within the dataset, our analysis is confined to the final, scan-converted images. Similar approach might be implemented in raw data as well. In our endeavor to facilitate outpainting, we meticulously segmented the cone from the image data through the application of conventional image processing techniques. Initially, we located the first non-zero pixel from the top as a means to identify the tip of the probe. Subsequently, we delineated the furthest left and right boundaries emanating from the probe to the edge of the image. By determining the angular relationship between these boundaries, we systematically removed segments measuring 15, 23, 30, and 40 degrees from each side in a symmetrical fashion. The resulting masks, corresponding to the excised segments, were preserved for further analysis.
echoGAN architecture
The proposed GAN architecture consists of two neural networks : generator and discriminator.
  1. a)
@@ Methods
    As a generator we employed U-Net architecture, as introduced by Ronneberger et al. [16]. It features two input channels: one for the masked image and the other for the binary mask, which indicates the regions for outpainting. It is configured with four downsampling layers, each characterized by specific kernel sizes [7, 5, 5, 5] and strides [1, 2, 2, 2], optimizing the network’s capacity for feature extraction and image reconstruction within the designated areas.
  2. b)
@@
Experimental design
@@
where $\mu$ is mean and $\sum$ is covariance of of the feature vectors for the real and generated images respectively. $Tr$ is the trace. A lower FID indicates better similarity between the distributions, suggesting higher image quality and fidelity.
Statistical analysis
@@
To statistically evaluate the difference between the real and AI-generated right ventricle area measurements, we first assessed whether each group of measurements conformed to a normal distribution. This preliminary step involved applying the Shapiro-Wilk test [18], a widely accepted method for testing normality. The results of the Shapiro-Wilk test indicated that the data do not follow a normal distribution, thereby necessitating the selection of a non-parametric statistical test for further analysis. Consequently, we opted for the Wilcoxon signed-rank test, a non-parametric counterpart to the paired t-test, which is designed for comparing two related samples. This test was employed to examine our hypothesis that there is no significant statistical difference between the areas determined from the AI-generated and real echocardiography data. The choice of the Wilcoxon signed-rank test was appropriate given the non-normal distribution of our data, allowing us to conduct a robust comparison of the median values between the two groups without the assumption of normality. The statistical analysis was conducted using the Python programming language, utilizing the SciPy library.
Acknowledgements (not compulsory)
Author contributions statement
Must include all authors, identified by initials, for example: A.A. conceived the experiment(s), A.A. and B.A. conducted the experiment(s), C.A. and D.A. analysed the results. All authors reviewed the manuscript.
Additional information
To include, in this order: Accession codes (where applicable); Competing interests (mandatory statement).
The corresponding author is responsible for submitting a competing interests statement on behalf of all authors of the paper. This statement must be included in the submitted article file.

### v2

# echoGAN: Extending the Field of View in Transthoracic Echocardiography Through Conditional GAN-Based Outpainting

@@ Abstract
Transthoracic Echocardiography (TTE) is a fundamental, non-invasive diagnostic tool in cardiovascular medicine, enabling detailed visualization of cardiac structures crucial for diagnosing various heart conditions. Despite its widespread use, TTE ultrasound imaging faces inherent limitations, notably the trade-off between field of view (FoV) and resolution. This paper introduces a novel application of conditional Generative Adversarial Networks (cGANs), specifically designed to extend the FoV in TTE ultrasound imaging while maintaining high resolution. Our proposed cGAN architecture, termed echoGAN, demonstrates the capability to generate realistic anatomical structures through outpainting, effectively broadening the viewable area in medical imaging. This advancement has the potential to enhance both automatic and manual ultrasound navigation, offering a more comprehensive view that could significantly reduce the learning curve associated with ultrasound imaging and aid in more accurate diagnoses. The results confirm that echoGAN reliably reproduce detailed cardiac features, thereby promising a significant step forward in the field of non-invasive cardiac naviagation.
keywords:
medical imaging , ultrasound
journal: Computers in Biology and Medicine
[…]
[TUKE]organization=Technical University of Kosice, Slovakia, addressline=Letna 9, city=Kosice, postcode=040 01, country=Slovakia
[…]
[POLYMTL]organization=Polytechnique Montréal, addressline=2500, chemin de Polytechnique, city=Montreal, postcode=H3T 1J4, country=Canada
[…]
[CORTEXVISION]organization=CortexVision, ddressline=Turgenevova 33 , city=Košice, postcode=040 01, country=Slovakia
1 Introduction
Transthoracic Echocardiography (TTE) Ultrasound (USG) is a cornerstone diagnostic tool in cardiovascular medicine, valued for its non-invasive approach and ability to comprehensively assess cardiac anatomy and function [1, 2]. It enables detailed visualization of cardiac structures, such as chambers, valves, and myocardium, and is essential in diagnosing conditions like heart failure, valvular heart disease, and cardiomyopathies. One of TTE USG’s key advantages is its capacity for real-time dynamic imaging, allowing clinicians to observe cardiac motion and provide insights beyond what static imaging can offer.
@@ 1 Introduction
In this context, our research has explored the application of a conditional Generative Adversarial Networks (cGANs) to extend the FoV in TTE USG imaging while maintaining high resolution. The proposed cGAN architecture, denoted as echoGAN, is capable of generating realistic anatomical structures through outpainting, thereby broadening the viewable area in medical imaging. This approach aims to address the limitations in FoV and resolution trade-offs, potentially enhancing both automatic and manual ultrasound navigation. For automatic navigation [3], an expanded FoV can provide a more comprehensive view, potentially improving the efficiency and accuracy of imaging processes, especially in complex diagnostic scenarios. Similar challenges in expanding the FoV have been addressed in related fields, using advanced lens models and flexible array technology to enhance the visual scope and resolution [4, 5].
@@
Furthermore, this extended FoV could be particularly beneficial for less experienced physicians in manual navigation.The study by Chisholm et al. [6] and the work of Vinodth et al. [7] focus on cardiac ultrasound training, specifically evaluating the duration of training necessary to achieve proficiency in TTE. These studies demonstrate that mastering TTE requires a considerable amount of time. By offering a wider yet detailed view, our cGAN-based approach could reduce the learning curve associated with ultrasound imaging. This process is often not straightforward and demands a high level of dexterity and spatial awareness. Given these challenges, there is a continuous search for technological advancements that can aid in simplifying the navigation process, making echocardiography more accessible and reliable, especially for less experienced practitioners. Moreover, the utilization of semi-synthetic data to enhance training models for medical device segmentation [8] underscores the potential for augmented datasets to refine AI-driven tools.
@@
In the domain of medical imaging, particularly ultrasound imaging, the task of outpainting introduces a set of unique challenges that are markedly distinct from those encountered in the outpainting of general or natural images [10, 11, 12]. The paramount requirement for anatomical accuracy in medical images necessitates that outpainting algorithms not only generate visually plausible extensions but also ensure these extensions accurately represent the expected anatomy. This is crucial, as any deviation could compromise the clinical utility of the image. Moreover, ultrasound images are characterized by specific textures and patterns that denote various tissues, fluids, and pathological conditions.
The paper is organized as follows. In the subsequent section we provide description of the proposed approach, detailing echoGan framework which leverages domain-aware augmentation and conditional GAN architecture. The third section describes the datasets utilized to validate the proposed approach. The fourth section discusses the results obtained from our experiments. Following discussion, the paper concludes with final remarks on the study’s findings and implications.
2 Field of View Outpainting
@@
This approach not only promises to refine navigation accuracy but also to facilitate a more comprehensive understanding of the heart’s structure and function. Enhancing the FoV through outpainting reduces navigation difficulty and provides a more comprehensive and detailed view. The proposed approach aims to address longstanding challenge of incomplete cardiac visualization during echocardiography.
2.1 Problem formulation
Given an input image $x\in\mathbb{R}^{H\times W}$, where $H$ and $W$ represent the height and width of the original image, and a binary mask $m\in\{0,1\}^{H\times W}$ that defines the region to be outpainted (with $1$ indicating pixels to be retained and 00 indicating the masked region), the task is to generate an extended output image $\hat{y}\in\mathbb{R}^{H\times(W+\Delta W)}$ that fills in the missing content in the masked area while preserving visual and anatomical coherence.
The problem can be expressed as learning a mapping function $G:(\mathbb{R}^{H\times W}\times\{0,1\}^{H\times W})\rightarrow\mathbb{R}^{H\times(W+\Delta W)}$ such that:
$\hat{y}=G(x\odot(1-m),m)$ (1)
where $\odot$ denotes the element-wise multiplication operator and $x\odot(1-m)$ is the input image with the masked region removed, leaving only the known parts.
The objective is to train $G$ such that the generated image $\hat{y}$ is indistinguishable from the true extended image $y$ under a discriminator $D$.
2.2 echoGAN
2.2.1 Domain-aware augmentation
We propose a novel domain-aware augmentation technique specifically designed for ultrasound imaging. This augmentation simulates the cone-shaped FOV typical in ultrasound by masking the side regions of the cone. The augmented mask enforces the generator to reconstruct the occluded areas.
For ultrasound image $x$, the cone-like FOV is defined by a binary mask $m_{\text{cone}}\in\{0,1\}^{H\times W}$, where
$m_{\text{cone}}(i,j)=\begin{cases}1,&\text{if }\tan^{-1}\left(\frac{|j-c_{x}|}{H-i}\right)\leq\frac{\theta}{2},\\ 0,&\text{otherwise}.\end{cases}$ (2)
Here, $m_{\text{cone}}(i,j)=1$ for pixels inside the cone-like region and $m_{\text{cone}}(i,j)=0$ for those outside the cone (cut-out regions). The cone is parameterized by the center of the cone $(c_{x},c_{y})$ and the angular spread $\theta$.
The augmentation simulates missing information by masking the side regions of the cone, requiring the generator to infer and reconstruct the missing areas by utilising augmented mask $m_{\text{augment}}\in\{0,1\}^{H\times W}$ defined as
$m_{\text{augment}}(i,j)=\begin{cases}1,&\text{if }\tan^{-1}\left(\frac{|j-c_{x}|}{H-i}\right)\leq\frac{\theta_{\text{shrink}}}{2},\\ 0,&\text{otherwise}.\end{cases}$ (3)
where $\theta_{\text{shrink}}<\theta$ controls the extent of the occlusion on the sides.
Then, augmented input image is defined as
$z_{\text{aug}}=\{x\odot m_{\text{augment}},m_{\text{augment}}\},$ (4)
where $x\odot m_{\text{augment}}$ represents ultrasound masked image retaining only the visible region.
Finally, the generator $G$ takes the augmented input $z_{\text{aug}}$ and produces the extended image $\hat{y}$:
$\hat{y}=G(z_{\text{aug}}).$ (5)
2.2.2 Network architecture and training
@@
The proposed GAN architecture consists of two neural networks : generator and discriminator. As a generator we employed U-Net architecture, as introduced by Ronneberger et al. [13]. It features two input channels: one for the masked image and the other for the binary mask, which indicates the regions for outpainting. It is configured with four downsampling layers, each characterized by specific kernel sizes [7, 5, 5, 5] and strides [1, 2, 2, 2], optimizing the network’s capacity for feature extraction and image reconstruction within the designated areas.
@@
The objective is to find the optimal generator $G^{*}$ that minimizes the combined loss $\mathcal{L}$ while simultaneously training the discriminator $D$ to maximize $\mathcal{L}_{D}$:
$G^{*}=\arg\min_{G}\max_{D}\left[\mathcal{L}_{D}+\mathcal{L}\right].$ (10)
This formulation ensures that the generator produces realistic outpainted images that closely resemble true TTE ultrasound images.
3 Data
We utilised two datasets in this paper: CAMUS dataset [14] collected under the CAMUS project [14] at the University Hospital of St. Étienne and EchoNet Dynamic echocardiography dataset [15].
CAMUS dataset was designed to facilitate measurements of Left Ventricular Ejection Fraction (LVEF). For each patient, 2D apical four-chamber and two-chamber view sequences were recorded. These sequences were extracted from video recordings, yielding an average of 20 images per patient (minimum: 10, maximum: 42). In total, the dataset comprises 9,964 images, of which 7,917 images from 400 patients were allocated for training a generative adversarial network. To prevent data leakage, all data from a given patient were confined to either the training set or the testing set, ensuring no overlap.
To further validate the echoGAN, we incorporated also the EchoNet Dynamic, consisting of 10,030 apical four-chamber echocardiography videos obtained from routine clinical care at Stanford University Hospital between 2016 and 2018. Each video provides a visualization of the heart from various angles and positions using different image acquisition techniques. To standardize the dataset, all videos were cropped and masked to exclude textual information and irrelevant content outside the scanning sector. The resulting frames were subsequently downsampled using cubic interpolation to a uniform resolution of 112 × 112 pixels.
Due to the absence of US data in raw format within the datasets, our analysis is confined to the final, scan-converted images. A similar approach might be implemented in raw data as well. We segmented the cone from the image data through the application of conventional image processing techniques. Initially, we located the first non-zero pixel from the top as a means to identify the tip of the probe. Subsequently, we delineated the furthest left and right boundaries emanating from the probe to the edge of the image. By determining the angular relationship between these boundaries, we systematically removed segments measuring 15, 23, 30, and 40 degrees from each side in a symmetrical fashion. The resulting masks, corresponding to the excised segments, were preserved for further analysis.
4 Results
@@
where $\mu$ is mean and $\sum$ is covariance of of the feature vectors for the real and generated images respectively [16]. $Tr$ is the trace. A lower FID indicates better similarity between the distributions, suggesting higher image quality and fidelity.
4.1 Different size of extended view
We evaluated the influence of the training cone size on the quality of outpainted sections in ultrasound images. Figure 4 illustrates two distinct scenarios: the first utilizes a narrower 10-degree cone cut as the inference input, with the echoGAN algorithm tasked with generating outpainted extensions that span 40 degrees on both sides; the second employs a 60-degree cone cut for inference, with outpainting applied to 15-degree segments on each side. Notably, the generated images appear highly realistic, even when echoGAN is interfaced with only a 10-degree cut. The algorithm demonstrates the ability to reproduce cardiac muscle structures with high fidelity. This is particularly evident in the first, second, and fourth rows, where echoGAN successfully outpaints a complete right atrium; that is an encouraging result given that the right atrium was not included in the 10-degree cut, suggesting that echoGAN can reconstruct it entirely. In the case of the left ventricle and left atrium, we observe that echoGAN is capable of outpainting the missing boundary parts for the 30-degree cut and can significantly reconstruct substantial portions of these structures for the 10-degree cut. Again, to quantitatively evaluate the quality of outpainted USG images, we measure FID for different sizes of outpainted regions. The results are provided in Tab. 1. The 30-degree cut means that echoGAN generates 15 degrees on both sides. Similarly, in the case of 80 degrees, echoGAN generates 40 degrees on both sides, using only a very narrow 10-degree cone as an input.
Method Metric CAMUS EchoNet Dynamic
Cut 30 Cut 46 Cut 60 Cut 80 Cut 30 Cut 46 Cut 60 Cut 80
echoGAN MSE $\downarrow$ 0.004893 0.008215 0.00173 0.00124 0.01057 0.01064 0.0175
L1 $\downarrow$ 0.039 0.05033 0.063 0.06114 0.05719 0.05653 0.08762
FID $\downarrow$ 122.08 143.83 145.74 149.63
LPIPS $\downarrow$ 0.05304 0.09 0.126 0.1211 0.1583 0.1567 0.2174
PSNR $\uparrow$ 23.91 21.46 19.77 19.95 20.15 20.13 17.89
SSIM $\downarrow$ 0.2749 0.3088 0.4707 0.4596 0.4799 0.4815 0.5881
CoMoDoGAN [17] MSE $\downarrow$ 0.1376 0.16 0.178 0.1720
L1 $\downarrow$ 0.32 0.368 0.4 0.3902
FID $\downarrow$ 109.59 158.68 208.59 194.03
LPIPS $\downarrow$ 0.1642 0.2068 0.2732 0.2584
PSNR $\uparrow$ 11.589 9.3082 7.5371 7.7864
SSIM $\downarrow$ 0.5674 0.7043 0.7973 0.7974
Table 1: The results for different sizes of outpainted area across CAMUS and EchoNet datasets.
4.2 Right ventricle area
Our experiments are specifically aimed at the quantification of the RV area from both original and outpainted echocardiograms. The rationale for focusing on the RV area is influenced by the established practice in cardiology where, for instance, the measurement of left ventricular ejection fraction incorporates the left ventricular area as a crucial parameter in its evaluation [19]. Drawing inspiration from such methodologies, we hypothesize that if the RV areas derived from generated (outpainted) and real echocardiograms are similar, it would substantiate the utility and validity of the generated images for clinical purposes.
@@ 4.2 Right ventricle area
To statistically evaluate the difference between the real and AI-generated right ventricle area measurements, we first assessed whether each group of measurements conformed to a normal distribution. This preliminary step involved applying the Shapiro-Wilk test [20], a widely accepted method for testing normality. The results of the Shapiro-Wilk test indicated that the data do not follow a normal distribution, thereby necessitating the selection of a non-parametric statistical test for further analysis. Consequently, we opted for the permutation test, a robust non-parametric approach that does not rely on assumptions of normality or symmetry. This test was employed to examine our hypothesis that there is no significant statistical difference between the areas determined from the AI-generated and real echocardiography data. The choice of the permutation test was appropriate given the non-normal distribution of our data, allowing us to conduct a reliable comparison of the paired differences between the two groups. The results of this test confirmed the null hypothesis ($T=8.804$, $p=0.762$), indicating that there is no statistically significant difference between the RV areas measured in the original echocardiographic images and those derived from images extended through echoGAN outpainting. This finding validates the accuracy and effectiveness of the GAN-based outpainting technique in echocardiography. The ability of GANs to generate anatomically coherent extensions of the cardiac structures without compromising the integrity of the images suggests that this approach can reliably extend the field of view in echocardiographic navigation. The statistical analysis was conducted using the Python programming language, utilizing the SciPy library.
Figure 5: Segmentation over Right Ventricle for GT 4CH view
5 Discussion
6 Conclusion
This paper introduced echoGAN, an innovative approach leveraging conditional Generative Adversarial Networks (cGANs) to extend the field of view (FoV) in Transthoracic Echocardiography (TTE) ultrasound imaging. By addressing the inherent trade-offs between FoV and resolution in standard ultrasound imaging, echoGAN is capable of producing realistic outpainted regions, enhancing the visualization of cardiac structures while maintaining high-resolution detail. This capability can significantly aid both manual and automated ultrasound navigation, reducing the learning curve for less experienced operators and facilitating more accurate diagnoses.
Quantitative analysis demonstrated the robustness of echoGAN, with a Fréchet Inception Distance (FID) of 122.08 for a 30-degree outpainting scenario and slightly higher FID values for larger outpainted regions, such as 143.83 for a 46-degree cut and 149.63 for an 80-degree cut. These results indicate that echoGAN effectively maintains anatomical coherence and quality even as the scope of outpainting increases. Clinical validation through the comparison of right ventricular (RV) area measurements in original and outpainted images further reinforced the reliability of this approach, with no statistically significant differences found, as verified by the Wilcoxon signed-rank test.
While echoGAN shows significant potential, limitations include the need for more diverse datasets and improvements in generating artifact-free extensions, particularly in complex anatomical areas. Future work should explore integrating shape priors from CT or MRI scans and include non-standard cardiac views to enhance anatomical fidelity.
Acknowledgment
Funded by the EU NextGenerationEU through the Recovery and Resilience Plan for Slovakia under the project No. 09I03-03-V04-00394 and project No. 09I03-03-V04-00639.
(Part of the) Research results was obtained using the computational resources procured in the national project National competence centre for high performance computing (project code: 311070AKF2) funded by European Regional Development Fund, EU Structural Funds Informatization of society, Operational Program Integrated Infrastructure